OpenStudy (anonymous):

The slope of the line containing the points (-5, 3) and (-2, 1) is a.-2/3 b.2/3 c.-3/2

OpenStudy (anonymous):

\[\frac{ 1-3 }{-2--5} =?\]

OpenStudy (anonymous):

Do you get how to figure it out from here?

OpenStudy (anonymous):

no :/

OpenStudy (jhannybean):

take (x1 , y1) = (-5 , 3) (x2 , y2) = (-2 , 1) (y2 - y1) /(x2 - x1) = (1 - 3) / (-2 - (-5)) simplify

OpenStudy (jhannybean):

You have - 2 + 5 , what does that give you?

OpenStudy (anonymous):

Okay, so subtract 3 from 1. You get -2. Then, add -2 and 5. The answer is \[-\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

thanks what about this one The slope of a line that is perpendicular to a line whose slope is 3/8 is a.-8/3 b.3/8 c.8/3

OpenStudy (jhannybean):

a perpendicular line has a negative and opposite slope original line slope = m perpendicular line = - 1/ m Can you go from there? :)

OpenStudy (anonymous):

no sorry

OpenStudy (jhannybean):

m = 3/8 perpendicular = -1 / (3/8) = -1/3 * 8/1 simplify this and you'll have your answer.

OpenStudy (anonymous):

so the answer is 8/3 right??

OpenStudy (jhannybean):

you forgot to multiply by the negative :)

OpenStudy (anonymous):

oh so -8/3

OpenStudy (anonymous):

Correct :)